OpenStudy (anonymous):

Help help help help please (metal) using pythagorean inequalities, determine which of the following sets of values can represent the lengths of the sides of an acute triangle. A. 1,2,3 B. 6,8,10 C. 6,7,8 D. 4,6,10

OpenStudy (anonymous):

@kingdeolt

OpenStudy (anonymous):

AB = [5,6,7] length sqrt(110) AC = [5,5,5] length sqrt(75) CB = [0,1,2] length sqrt (5) So pythagorean inequality states that for a<b<c if a^2 + b^2 < c^2, then the triangle is acute. here CB < AC < AB So try doing (CB)^2 + (AC)^2 = 75 + 5 = 80 which is less than (AB)^2 being 110. Is this what you are looking for?